OpenStudy (anonymous):

How many more candies could machine J pack than machine K in 11 minutes? 30 44 90 330

OpenStudy (anonymous):

http://prntscr.com/6zp7wa @Nnesha

OpenStudy (ineedhelpfast12):

30

Nnesha (nnesha):

x represent what ?

OpenStudy (anonymous):

the minutes

OpenStudy (ineedhelpfast12):

26x represents the slope. you have to solve for the y intercept, which ends up being 30

Nnesha (nnesha):

okay so x =11 replace x by 11 and then solve for y ^^^ that's for machine K

OpenStudy (anonymous):

Wait I'm confused where'd you get 11 from?

OpenStudy (ineedhelpfast12):

Hey do you get how I got 30?

OpenStudy (anonymous):

Not really

Nnesha (nnesha):

11 is in the question

OpenStudy (anonymous):

Oh yeah, sorry.

OpenStudy (ineedhelpfast12):

ok. so i found the amount it can pack in 1 minute, so multiply 30 by 11

OpenStudy (anonymous):

330

OpenStudy (ineedhelpfast12):

then you get how much more J can pack than K in 11 minutes

OpenStudy (ineedhelpfast12):

exactly @dirtydan69

OpenStudy (anonymous):

330 is what they get in j right?

OpenStudy (ineedhelpfast12):

yup

OpenStudy (anonymous):

so now I need to find what they get in K

OpenStudy (ineedhelpfast12):

now you subtract 26 times 11 from 330

OpenStudy (ineedhelpfast12):

the difference, or answer to this subtraction problem would be what?

OpenStudy (anonymous):

26 * 11 = 286 330-286= 44

OpenStudy (ineedhelpfast12):

good job! J can pack 44 more candies than K in 4 minutes

OpenStudy (ineedhelpfast12):

11, not 4 my bad!

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (ineedhelpfast12):

no problem. I shalt give thee a medal

OpenStudy (ineedhelpfast12):

sure

OpenStudy (anonymous):

One triangle on a graph has a vertical side of 5 and a horizontal side of 11. Another triangle on a graph has a vertical side of 15 and a horizontal side of 33. Could the hypotenuses of these two triangles lie along the same line? No, because they need to be the same size No, because they are not similar triangles Yes, because all triangles can fit along this line Yes, because they are similar triangles

OpenStudy (anonymous):

5*3=15 11*3=33 so they're different sizes but they're similar

OpenStudy (ineedhelpfast12):

yes because they are similar triangles

OpenStudy (anonymous):

so d is right??

OpenStudy (ineedhelpfast12):

yerp

OpenStudy (anonymous):

thanks so much

OpenStudy (ineedhelpfast12):

no problem

OpenStudy (ineedhelpfast12):

any more?

OpenStudy (anonymous):

no more

OpenStudy (ineedhelpfast12):

kk

OpenStudy (ineedhelpfast12):

later aligator